2. If the ultimate shear strength of a steel plate is 35000psi, what force is

necessary to punch a 0.35 in. diameter hole in a 0.255 in. thick plate?


( Ans F=9813.55 lb )


1
Expert's answer
2022-03-31T03:32:02-0400
"\sigma=\frac{F}{A}"

"F=\sigma A=\sigma\pi d t"

"F={\rm 35000\: psi\times3.14159\times 0.35\: in\times 0.255\: in}\\
=9813.55\:\rm lb"